# FancyVideo: Towards Dynamic and Consistent Video Generation via Cross-frame Textual Guidance

### v2

@@ Introduction
Text-Video Alignment CLIP-Socre($\uparrow$) 20.46 20.53 20.66 20.36 20.80 19.70 20.85
@@
Detection-Score($\uparrow$) 58.99 64.05 58.63 50.01 64.04 51.19 66.66
Color-Score($\uparrow$) 34.35 37.56 48.55 38.72 45.65 42.39 51.09
@@
Motion Quality Action Score($\uparrow$) 71.81 62.53 81.56 72.12 72.22 61.94 72.99
Motion AC-Score($\rightarrow$) 44 44 50 42 46 32 52
Flow-Score($\rightarrow$) 0.50 0.70 2.07 6.99 0.96 2.403 1.7413
@@
Temporal Consistency CLIP-Temp($\uparrow$) 99.97 99.94 99.77 99.74 99.75 99.85 99.84
Warping Error($\downarrow$) 0.0006 0.0008 0.0067 0.0162 0.0054 0.0177 0.0051
Face Consistency($\uparrow$) 99.62 99.06 99.32 98.94 99.34 99.63 99.31
@@
Experiments
In the quantitative experiments, FancyVideo utilizes the T2I base model to generate images as the first frame. In the qualitative experiments, for aesthetic purposes and to remove watermarks, an external model is used to generate a beautiful first frame. Experimental setup about training data, settings, and hardware is provided in the supplementary material.
Qualitative Evaluation
We choose AnimateDiff (Guo et al. 2023b), DynamiCrafter (Xing et al. 2023), and two commercialized products, Pika (PikaLabs 2024) and Gen2 (Runway 2024), for a composite qualitative analysis. It is worth noting that in the quantitative experiments, the first frame of FancyVideo is generated by SDXL to achieve a more aesthetically pleasing result and to minimize the appearance of watermark (although subsequent frames may still exhibit it).
As shown in Fig. 4, our approach exhibits superior performance, outperforming previous methods regarding temporal consistency and motion richness. In contrast, AnimateDiff, DynamiCrafter, and Gen2 generate videos with less motion. Pika struggles to produce object-consistent and high-quality video frames. Remarkably, our method can accurately understand the motion instructions in the text prompt (e.g., ”A teddy bear walking … beautiful sunset.” and ”A teddy bear running … City.” case). More videos and applications (e.g., Personalized Video Generation and Video Extending) are provided in our supplementary material.
Quantitative Evaluation
For a comprehensive comparison with the SOTA methods, we adopt three popular benchmarks (e.g., EvalCrafter (Liu et al. 2023), UCF-101 (Soomro, Zamir, and Shah 2012), and MSR-VTT (Xu et al. 2016) ) and human evaluation to evaluate the quality of video generation. Among them, EvalCrafter is a relatively comprehensive benchmark for video generation currently. UCF-101 and MSR-VTT are benchmarks commonly used in previous methods (Girdhar et al. 2023; Zhang et al. 2023b). Meanwhile, human evaluation can compensate for the inaccuracies in existing text-conditioned video generation evaluation systems.
@@
EvalCrafter (Liu et al. 2023) quantitatively evaluates the quality of text-to-video generation from four aspects (including Video Quality, Text-video Alignment, Motion Quality, and Temporal Consistency). Each dimension contains multiple subcategories of indicators shown in the Table. 1. As discussed in community (Liu and Cun 2024), the authors acknowledge that the original manner of calculating the comprehensive metric was inappropriate. For a more intuitive comparison, we introduce a comprehensive metric for every aspect by considering each sub-indicators numerical scale and positive-negative attributes. Detailed information about the comprehensive metric can be available in supplementary material.
In detail, we compare the performance of the previous video generation SOTA methods (e.g., Pika (PikaLabs 2024), Gen2 (Runway 2024), Show-1 (Zhang et al. 2023b), ModelScope (Wang et al. 2023a), DynamiCrafter (Xing et al. 2023), and AnimateDiff (Guo et al. 2023b)) and exhibit in Table. 1. Our method demonstrates outstanding performance beyond existing methods at the Video Quality and Text-video Alignment aspect. Although Show-1 has the best Motion Quality (81.56), its Video Quality is poor (only 85.08). That indicates that it cannot generate high-quality videos with reasonable motion. However, our method has the second highest Motion Quality (72.99) and the best Video Quality (177.72), achieving the trade-off between quality and motion. The above results indicate the superiority of FancyVideo and its ability to generate temporal-consistent and motion-accurate video.
Table 2: Quantitative evaluation on the UCF-101 (Soomro, Zamir, and Shah 2012) and MSR-VTT (Xu et al. 2016) . The best and second performing metrics are highlighted in bold and underline respectively.
@@
Emu Video (2023) 34M 606.20 42.70 - - -
AnimateDiff (2023b) 10M 584.85 37.01 61.24 628.57 0.2881
DynamiCrafter (2023) 10M 404.50 41.97 32.35 219.31 0.2659
Show-1 (2023b) 10M 394.46 35.42 - 538.00 0.3072
ModelScope (2023a) 10M 410.00 - - 550.00 0.2939
@@
Following the prior work (Zhang et al. 2023b), we evaluate the zero-shot generation performance on UCF-101 (Soomro, Zamir, and Shah 2012) and MSR-VTT (Xu et al. 2016) as shown in Table. 2. We use Fréchet Video Distance (FVD) (Unterthiner et al. 2019), Inception Score (IS) (Wu et al. 2021), Fréchet Inception Distance (FID) (Heusel et al. 2017), and CLIP similarity (CLIPSIM) as evaluation metrics and compared some current SOTA methods. FancyVideo achieves competitive results, particularly excelling in IS and CLIPSIM with scores of 43.66 and 0.3076, respectively. Besides, previous studies (Ho et al. 2022; Girdhar et al. 2023; Wu et al. 2023b) have pointed out that these metrics do not accurately reflect human perception and are affected by the gap between the distribution of training and test data and the image’s low-level detail.
Table 3: Ablation studies on the core component of FancyVideo.
TAR TFB TII Video Quality Text-Video Alignment Motion Quality Temporal Consistency
@@
Human Evaluation.
Inspired by EvalCrafter (Liu et al. 2023), we introduce a multi-candidate ranking protocol with four aspects: video quality, text-video alignment, motion quality, and temporal consistency. In this protocol, participants rank the results of multiple candidate models for each aspect. Each candidate model receives a score based on its ranking. For instance, if there are $N$ candidate models ranked by video quality, the first model gets $N-1$ points, the second gets $N-2$ points, and so on, with the last model receiving 0 points. Adhering to this protocol, we selected 108 samples from the EvalCrafter validation set and gathered judgments from 100 individuals. As depicted in Fig. 5, our method significantly outperforms text-to-video conversion methods, including AnimateDiff (Guo et al. 2023b), Pika (PikaLabs 2024), and Gen2 (Runway 2024), across all four aspects. FancyVideo demonstrates exceptional motion quality while preserving superior text-video consistency. Additionally, we conducted a similar comparison of four image-to-video methods, including DynamiCrafter (Xing et al. 2023), Pika, and Gen2, as shown in Fig. 5.
Ablation Studies
In this section, we conduct the experiment and exhibit the visual comparison on the EvalCrafter (Liu et al. 2023) to explore the effect of critical design in CTGM. Specifically, the ablation component contains the three core modules (TII, TAR, and TFB). As shown in Table. 3, TAR can effectively improve the performance on both metrics, showing that temporal refined attention map operations are beneficial for video generation. Continued insertion of TFB and TII further improves the generator’s performance, due to the refined latent features and frame-level personalized text condition. Meanwhile, the qualitative analysis is included in the appendix.
Conclusion
@@
Appendix A Supplementary material
@@
  • (Sec. 2) More details on evaluation metrics employed in UCF101, MSR-VTT, human evaluation and the EvalCrafter (Liu et al. 2023).
@@
Experimental Setup
@@
We utilize WebVid-10M (Bain et al. 2021) as the training data. The WebVid-10M dataset contains 10.7 million video-caption pairs, with most videos having a resolution of 336 $\times$ 596. Since every clip in the WebVid-10M has a watermark, the generated video inevitably appears watermarked.
@@
FancyVideo is trained on the WebVid-10M dataset. The video clips are initially sampled with a stride of 4, followed by resizing and center-cropping to a resolution of 256 × 256. We utilize Stable-Diffusion v1.5 (Rombach et al. 2022) as the text-to-image (T2I) base model and train exclusively with the temporal attention block and our CTGM block. For the 16-frame training, we use a batch size of 512 and train for 12,000 iterations. For training with 32 frames, 48 frames, and 64 frames, we use batch sizes of 256, 256, and 128, respectively. The training process comprises 24,000 iterations for 32 frames, 48,000 iterations for 48 frames, and 96,000 iterations for 64 frames on 64 A100 GPUs with 80G memory. At inference, the sampling strategy for video generation is DDIM (Song, Meng, and Ermon 2020) with 50 steps. Also, we utilize the classifier-free guidance (Ho and Salimans 2022) with a 7.5 guidance scale. Similar to AnimateDiff (Guo et al. 2023b), we swap the base model with Realistic-Vision v5.1. The evaluations on EvalCrafter (Liu et al. 2023) and all the qualitative experiments are conducted at a resolution of 512. Regarding the evaluations on UCF-101 (Soomro, Zamir, and Shah 2012) and MSR-VTT (Xu et al. 2016), following (Zhang et al. 2023b), we conduct assessments on videos generated at a resolution of 256.
More Details on Evaluation Metrics
Details of evaluation on UCF101
@@
To calculate the Fréchet Video Distance (FVD) (Unterthiner et al. 2019) and Inception Score(IS) (Blattmann et al. 2023b; Ren et al. 2024; Saito et al. 2020), we produce 2048 videos based on the class distribution of the UCF101 dataset. These videos are generated at a resolution of 256 pixels. Subsequently, we extract I3D embeddings from our videos. Next, we compute the FVD score by comparing the I3D embeddings of our videos with those of the UCF101 videos. For computing the Inception Score (IS), the same set of generated videos was utilized to extract C3D embeddings.
Details of evaluation on MSR-VTT
@@
The MSR-VTT dataset is an open-domain video retrieval and captioning dataset with 10,000 videos, each having 20 captions. The standard splits include 6,513 training videos, 497 validation videos, and 2,990 test videos. For our experiments, we use the official test split and randomly select a text prompt for each video during evaluation. Using TorchMetrics, we compute our CLIPSIM (Wu et al. 2021) metrics with the CLIP-VIT-B/32 model. We calculate the CLIP similarity for all frames in the generated videos and report the averaged results.
Details of Human Evaluation
@@
We filter out 108 prompts from the 700 prompts in EvalCrafter for human evaluation. This subset of prompts contains more detailed descriptions. The prompt-108 we utilized is presented in A.
About EvalCrafter
@@
As mentioned in EvalCrafter Benchmark., there are unreasonable aspects in how EvalCrafter calculates Video Quality, Text-Video Alignment, Motion Quality, and Temporal Consistency. Therefore, we propose our calculation scheme. Specifically, we remove some neutral sub-metrics and those with significant differences in scale, and based on whether each sub-metric is positive or negative, we obtain a comprehensive score. For Video Quality, it can be represented as follows:
$Video\_Quality=VQAA+VQAT+IS$ (8)
@@
Figure 6: Evaluation results of varying-frames video generation. EC denotes the comprehensive metric on EvalCrafter(Liu et al. 2023) benchmark and HE indicates Human Evaluation.
@@
$Motion\_Quality=Action\_Score$ (9)
@@
$Temporal\_Consistency=CLIP\_Temp+Face\_Consistency$ (10)
@@
More Applications
Given the flexibility of our method in swapping T2I base models, we conduct experiments similar to prior work (Guo et al. 2023b), with different base models. Fig. 7 displays the results of our method using models downloaded from the Civitai(Civitai 2024) community.
@@
Due to the flexibility of our framework, we can easily combine our method with text-to-image super-resolution modules (Cheng et al. 2024). We conduct experiments to achieve multi-scale and high-resolution video generations, as depicted in Fig. 12.
@@
More Generation Results
@@
Figure 10: Qualitative comparisons of w CTGM and w/o CTGM in our method. In the solid red box on the left, unnatural stillness is apparent. Meanwhile, there is distortion in the red dashed box on the right. In the setting with CTGM, the above issues are resolved.
The qualitative analysis in Fig. 10 demonstrates the experimental results of training with and without CTGM. In the upper case, without CTGM, the text is not fully understood, causing the subject in the red box to be unnaturally still. In bottom case, there is a distortion in the red dashed circle. With CTGM, our model generates more dynamic and undistorted videos.
Figure 11: Qualitative analysis of our approach to long video generation (64 frames). More results are shown in A.
@@
A list of prompts for human evaluation
  1. 1.

### v4

@@ 4.2 Quantitative Evaluation
CLIP-Score($\uparrow$) 20.46 20.53 20.66 20.36 20.80 19.70 20.85
@@
Text-Video Detection-Score($\uparrow$) 58.99 64.05 58.63 50.01 64.04 51.19 66.66
Alignment Color-Score($\uparrow$) 34.35 37.56 48.55 38.72 45.65 42.39 51.09
@@
Action Score($\uparrow$) 71.81 62.53 81.56 72.12 72.22 61.94 72.99
Motion Motion AC-Score($\rightarrow$) 44 44 50 42 46 32 52
Quality Flow-Score($\rightarrow$) 0.50 0.70 2.07 6.99 0.96 2.403 1.7413
@@
CLIP-Temp($\uparrow$) 99.97 99.94 99.77 99.74 99.75 99.85 99.84
Temporal Warping Error($\downarrow$) 0.0006 0.0008 0.0067 0.0162 0.0054 0.0177 0.0051
Consistency Face Consistency($\uparrow$) 99.62 99.06 99.32 98.94 99.34 99.63 99.31
@@
EvalCrafter Liu et al. (2023) quantitatively evaluates the quality of text-to-video generation from four aspects (including Video Quality, Text-video Alignment, Motion Quality, and Temporal Consistency). Each dimension contains multiple subcategories of indicators shown in the Table. 1. As discussed in community Liu and Cun (2024), the authors acknowledge that the original manner of calculating the comprehensive metric was inappropriate. For a more intuitive comparison, we introduce a comprehensive metric for every aspect by considering each sub-indicators numerical scale and positive-negative attributes.
In detail, we compare the performance of the previous video generation SOTA methods (e.g., Pika PikaLabs (2024), Gen2 Runway (2024), Show-1 Zhang et al. (2023), Lumiere Bar-Tal et al. (2024), DynamiCrafter Xing et al. (2023), and AnimateDiff Guo et al. (2023b)) and exhibit in Table. 1. Our method demonstrates outstanding performance beyond existing methods at the Video Quality and Text-video Alignment aspect. Although Show-1 has the best Motion Quality (81.56), its Video Quality is poor (only 85.08). That indicates that it cannot generate high-quality videos with reasonable motion. However, our method has the second highest Motion Quality (72.99) and the best Video Quality (177.72), achieving the trade-off between quality and motion. The above results indicate the superiority of FancyVideo and its ability to generate temporal-consistent and motion-accurate video.
@@
Emu Video 34M 606.20 42.70 - - -
AnimateDiff 10M 584.85 37.01 61.24 628.57 0.2881
DynamiCrafter 10M 404.50 41.97 32.35 219.31 0.2659
Show-1 10M 394.46 35.42 - 538.00 0.3072
Lumiere 10M 332.49 37.54 - 550.00 0.2939
@@
Table 2: Quantitative evaluation on the UCF-101 Soomro et al. (2012) and MSR-VTT Xu et al. (2016) .The best and second performing metrics are highlighted in bold and underline respectively.
@@
Following the prior work Zhang et al. (2023), we evaluate the zero-shot generation performance on UCF-101 Soomro et al. (2012) and MSR-VTT Xu et al. (2016) as shown in Table. 2. We use Fréchet Video Distance (FVD) Unterthiner et al. (2019), Inception Score (IS) Wu et al. (2021), Fréchet Inception Distance (FID) Heusel et al. (2017), and CLIP similarity (CLIPSIM) as evaluation metrics and compared some current SOTA methods. FancyVideo achieves competitive results, particularly excelling in IS and CLIPSIM with scores of 43.66 and 0.3076, respectively. Besides, previous studies Ho et al. (2022); Girdhar et al. (2023); Wu et al. (2023b) have pointed out that these metrics do not accurately reflect human perception and are affected by the gap between the distribution of training and test data and the image’s low-level detail.
Human Evaluation.
Inspired by EvalCrafter Liu et al. (2023), we introduce a multi-candidate ranking protocol with four aspects: video quality, text-video alignment, motion quality, and temporal consistency. In this protocol, participants rank the results of multiple candidate models for each aspect. Each candidate model receives a score based on its ranking. For instance, if there are $N$ candidate models ranked by video quality, the first model gets N1N-1 points, the second gets N2N-2 points, and so on, with the last model receiving 0 points. Adhering to this protocol, we selected 108 samples from the EvalCrafter validation set and gathered judgments from 100 individuals. As depicted in Fig. 5, our method significantly outperforms text-to-video conversion methods, including AnimateDiff Guo et al. (2023b), Pika PikaLabs (2024), and Gen2 Runway (2024), across all four aspects. FancyVideo demonstrates exceptional motion quality while preserving superior text-video consistency. Additionally, we conducted a similar comparison of four image-to-video methods, including DynamiCrafter Xing et al. (2023), Pika, and Gen2, as shown in Fig. 5.
4.3 Ablation Studies
In this section, we conduct extensive experiments and exhibit detailed visual comparisons on the EvalCrafter benchmark Liu et al. (2023) to thoroughly explore the effect of critical designs in CTGM. The ablation study includes two key modules (TII and TAR), each enhancing video quality As shown in Table 3, TAR significantly improves both metrics, highlighting the importance of temporal attention refinement. Adding TII further enhances performance by refining latent features and enabling frame-level text control.
TAR TII Video Text-Video Motion Temporal
Quality Alignment Quality Consistency
@@ 4.3 Ablation Studies
Table 3: Ablation studies on TAR and TII, where TFB is used by default in ablation experiments.
5 Conclusion
@@
Supplementary Material for
FancyVideo: Towards Dynamic and Consistent Video Generation via Cross-frame Textual Guidance
1 Appendix Section
@@
  • (Sec. 3) More details on evaluation metrics employed in UCF101, MSR-VTT, human evaluation and the EvalCrafter Liu et al. (2023).
@@
1.1 Temporal Feature Booster
The ablation study of the Temporal Feature Booster (TFB) is presented in Table 4.
TFB Video Text-Video Motion Temporal
Quality Alignment Quality Consistency
175.28 391.21 72.44 199.05
177.72 396.65 72.99 199.15
Table 4: Ablation studies on the TFB, where both TII and TAR are used by default in experimental configurations.
1.2 Experimental Setup
@@
We utilize WebVid-10M Bain et al. (2021) as the training data. The WebVid-10M dataset contains 10.7 million video-caption pairs, with most videos having a resolution of 336 $\times$ 596. Since every clip in the WebVid-10M has a watermark, the generated video inevitably appears watermarked.
@@
FancyVideo is trained on the WebVid-10M dataset. The video clips are initially sampled with a stride of 4, followed by resizing and center-cropping to a resolution of 256 × 256. We utilize Stable-Diffusion v1.5 Rombach et al. (2022) as the text-to-image (T2I) base model and train exclusively with the temporal attention block and our CTGM block. For the 16-frame training, we use a batch size of 512 and train for 12,000 iterations. For training with 32 frames, 48 frames, and 64 frames, we use batch sizes of 256, 256, and 128, respectively. The training process comprises 24,000 iterations for 32 frames, 48,000 iterations for 48 frames, and 96,000 iterations for 64 frames on 64 A100 GPUs with 80G memory. At inference, the sampling strategy for video generation is DDIM Song et al. (2020) with 50 steps. Also, we utilize the classifier-free guidance Ho and Salimans (2022) with a 7.5 guidance scale. Similar to AnimateDiff Guo et al. (2023b), we swap the base model with Realistic-Vision v5.1. The evaluations on EvalCrafter Liu et al. (2023) and all the qualitative experiments are conducted at a resolution of 512. Regarding the evaluations on UCF-101 Soomro et al. (2012) and MSR-VTT Xu et al. (2016), following Zhang et al. (2023), we conduct assessments on videos generated at a resolution of 256.
1.3 More Details on Evaluation Metrics
1.3.1 Details of evaluation on UCF101
@@
To calculate the Fréchet Video Distance (FVD) Unterthiner et al. (2019) and Inception Score(IS) Blattmann et al. (2023b); Ren et al. (2024); Saito et al. (2020), we produce 2048 videos based on the class distribution of the UCF101 dataset. These videos are generated at a resolution of 256 pixels. Subsequently, we extract I3D embeddings from our videos. Next, we compute the FVD score by comparing the I3D embeddings of our videos with those of the UCF101 videos. For computing the Inception Score (IS), the same set of generated videos was utilized to extract C3D embeddings.
1.3.2 Details of evaluation on MSR-VTT
@@
The MSR-VTT dataset is an open-domain video retrieval and captioning dataset with 10,000 videos, each having 20 captions. The standard splits include 6,513 training videos, 497 validation videos, and 2,990 test videos. For our experiments, we use the official test split and randomly select a text prompt for each video during evaluation. Using TorchMetrics, we compute our CLIPSIM Wu et al. (2021) metrics with the CLIP-VIT-B/32 model. We calculate the CLIP similarity for all frames in the generated videos and report the averaged results.
1.3.3 Details of Human Evaluation
@@
We filter out 108 prompts from the 700 prompts in EvalCrafter for human evaluation. This subset of prompts contains more detailed descriptions. The prompt-108 we utilized is presented in 1.6.
1.3.4 About EvalCrafter
@@
As mentioned in 4.2, there are unreasonable aspects in how EvalCrafter calculates Video Quality, Text-Video Alignment, Motion Quality, and Temporal Consistency. Therefore, we propose our calculation scheme. Specifically, we remove some neutral sub-metrics and those with significant differences in scale, and based on whether each sub-metric is positive or negative, we obtain a comprehensive score. For Video Quality, it can be represented as follows:
Video_Quality=VQAA+VQAT+ISVideo\_Quality=VQAA+VQAT+IS (8)
@@
Figure 6: Evaluation results of varying-frames video generation. EC denotes the comprehensive metric on EvalCrafter benchmark and HE indicates Human Evaluation.
@@
Motion_Quality=Action_ScoreMotion\_Quality=Action\_Score (9)
@@
Temporal_Consistency=CLIP_Temp+Face_ConsistencyTemporal\_Consistency=CLIP\_Temp+Face\_Consistency (10)
@@
1.4 More Applications
Given the flexibility of our method in swapping T2I base models, we conduct experiments similar to prior work Guo et al. (2023b), with different base models. Fig. 7 displays the results of our method using models downloaded from the CivitaiCivitai (2024) community.
@@ 1.4 More Applications
Due to the flexibility of our framework, we can easily combine our method with text-to-image super-resolution modules Cheng et al. (2024); Xu et al. (2025); Wang et al. (2025b, 2022). We conduct experiments to achieve multi-scale and high-resolution video generations, as depicted in Fig. 11.
@@
1.5 More Generation Results
@@
Figure 10: Qualitative analysis of our approach to long video generation (64 frames). More results are shown in 1.5.
@@
1.6 A list of prompts for human evaluation